OpenStudy (anonymous):

Assume that the fisherman can use between 3 and 6 lines. With 3 lines, the probability of a catch on each line is 0.77. With 4 this probability is 0.62, with 5 it is 0.51, and with 6 the probability is 0.42. How many lines should the fisherman use to maximize the probability of catching at least 2 fish?

OpenStudy (anonymous):

its not 4 or 5

OpenStudy (anonymous):

i got same, answer was 5 and it said it was incorrect.

OpenStudy (anonymous):

i got answer 3 with Pr=.4091

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Remember, in probability, each line will be treated as an individual event. The values calculated by @Calcmathlete are incorrect as the probability exceeds 1. Remember, for any given event the probability can NEVER be greater than 1 or less than 0. In all these scenarios, he has to catch AT LEAST 2 fish so there can be 2 or more fish in each case. So lets look at it case by case: with 3 lines: P (at least 2 fish) = P(3 fishes) or P(2 fishes) P (at least 2 fish) = (0.77^3) + [(0.77^2) * 0.23] P (at least 2 fish) = 0.5929 with 4 lines: P (at least 2 fish) = P(4 fishes) or P(3 fishes) or P(2 fishes) P (at least 2 fish) = (0.62^4) + [(0.62^3) * 0.38] + [0.62^2 + 0.38^2] P (at least 2 fish) = 0.2938 with 4 lines: P (at least 2 fish) = P(5 fishes) or P(4 fishes) or P(3 fishes) or P(2 fishes) P (at least 2 fish) = (0.51^5) + [(0.51^4) * 0.49)] + [(0.51^3) * (0.49^2)] + [0.51^2 + 0.49^3] P (at least 2 fish) = 0.1301 It is obvious from here that with 6 lines, this probability will decrease. So you use 3 lines. The probability will be 0.5929

OpenStudy (anonymous):

Oh. Sorry. I didn't know that.

OpenStudy (anonymous):

mubzz I sooo appreciate your help. I like Baye's Probability better haha it's easier to figure out

OpenStudy (anonymous):

Haha yea there are definitely several ways to approach any given problem. As long as you stick to the principles, no restriction on what method you use :D Good luck!

OpenStudy (anonymous):

Thank You!